OpenStudy (anonymous):

factor polynomial completely b^3-4b

OpenStudy (triciaal):

first divide by b what do you have

OpenStudy (triciaal):

I should say factor out b to get b(b^2 -4) do you recognize b^2 - 4 as the difference of 2 squares?

OpenStudy (anonymous):

b has a 1 in front of it but i dont know how to express them as a bionomial

OpenStudy (triciaal):

think about (a^2 - b^2) = (a + b)(a-b) and apply to your problem

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

so the answer would be (b-4)(b+4) ?

OpenStudy (triciaal):

the complete factorization b(b + 2)(b-2)

OpenStudy (triciaal):

if x^2 =4 then x = sqrt of 4 = +2/-2

OpenStudy (triciaal):

you forgot to take the sqrt

OpenStudy (anonymous):

oh ok thanks

OpenStudy (triciaal):

you are welcome